### v3

# DoubleDiffusion: Combining Heat Diffusion with Denoising Diffusion for Generative Learning on 3D Meshes

Abstract
This paper proposes DoubleDiffusion, a novel framework that combines heat dissipation diffusion and denoising diffusion for direct generative learning on 3D mesh surfaces. Our approach addresses the challenges of generating continuous signal distributions residing on a curve manifold surface. Unlike previous methods that rely on unrolling 3D meshes into 2D or adopting field representations, DoubleDiffusion leverages the Laplacian-Beltrami operator to process features respecting the mesh structure. This combination enables effective geometry-aware signal diffusion across the underlying geometry. As shown in Fig. 1, we demonstrate that DoubleDiffusion has the ability to generate RGB signal distributions on complex 3D mesh surfaces and achieves per-category shape-conditioned texture generation across different shape geometry. Our work contributes a new direction in diffusion-based generative modeling on 3D surfaces, with potential applications in the field of 3D asset generation. https://github.com/Wxyxixixi/DoubleDiffusion_3D_Mesh
Figure 1: We present DoubleDiffusion, a framework for generative learning on 3D manifold mesh. We use (a) the heat dissipation diffusion as the geometric feature extractor to efficiently aggregate vertice features over the spatial domain on mesh surfaces. We employ a (b) temporal denoising diffusion process to capture the distribution of the signal on mesh surfaces. Our model can effectively learn the RGB signal distribution defined on a specific manifold (e.g., (c) sinewave and (d) bunny), to generate different (c) handwritten digits [11] or (d) face looks [18]. Moreover, Our model can generate the texture on different manifold shape given different manifold surfaces from the (e) ShapeNetCore [7].
1 Introduction
Significant advancements have been made recently in generative learning using the diffusion probabilistic models [14, 33], particularly in domains involving images [29, 26], videos [15, 6], and 3D such as points clouds [23], triplanes [32], Nerf [21], mesh texture generation based on grid [22], etc. However, the extension of generative learning to 3D object surfaces, typically represented as curved surfaces or manifolds, remains under-explored due to the lack of effective learning strategies on these warped manifold surfaces. Several existing methods [8, 42] attempt to learn the texture distribution on 3D object manifold surfaces through unwarped 2D texture maps. While these approaches enable surface representation learning in 2D format and can leverage 2D image models to infer texture samples, they often suffer from the inconsistency along edges and fail to fully capture the intricate geometry of the 3D surface. Additionally, unwrapping a 3D surface onto a 2D texture map is ill-posed without a certain solution, thus different unwrapping techniques can significantly impact the final generative results.
While some methods used the point cloud as the manifold surface representation, the neighborhood relationship and the distance are computed through K-nearest neighbors (KNN), the distance are not ideally calculated along the manifold surface [27, 28]. Meanwhile, this approach lacks the precision needed for mesh-based structures, where each point’s neighboring relationships are defined by the mesh’s inherent geometry. Since the mesh is naturally a structured manifold, with the geodesic distance represented via the vertices connectivity, mesh present an ideal medium for diffusion-based generation directly on the surface.
In this work, we present a probabilistic diffusion framework on manifold, addressing the challenge of generative learning directly on the mesh surfaces. Our method captures the local geometric characteristics of 3D shapes without needing to flatten the surface to a 2D medium, ensuring consistency of the content generation across the manifold. Bronstein [5] points out that, Laplacian Beltrami (LB) operator can capture the differential geometric property on the manifold defining neighboring and distance relationship when the information interchanging among the points. To model the diffusion on the manifold mesh surface, we adapt from the mesh processing theory [43] and use the Laplacian Beltrami operator to capture the local geometric of the manifold mesh. Since mesh is a structured manifold, where the geometry of the manifold are explicitly entailed in the mesh topology, the geometric information of the manifold mesh, such as curvatures, geodesic distances of the mesh can be captured and process through the discrete Laplacian-Beltrami Operators [34, 43], opening up a great opportunity to directly learn a scalar distribution defined on the manifold domain.
Specifically, we employ the LB operator as a core feature processing element to approximate heat dissipation diffusion. This diffusion process models the principles of heat transfer on continuous surfaces, where heat spreads over time based on the geometry of the surface, allowing information to propagate across points in a structured manner [13, 9, 31]. On a mesh manifold, the LB operator can approximate this process, effectively simulating heat diffusion and enabling smooth information transfer across vertices. In this paper, we propose to integrate the heat dissipation diffusion with the diffusion probabilistic model [14] as the generative learning framework on the manifold. To our knowledge, we are the first work to combine heat dissipation diffusion with a denoising diffusion probabilistic model (DDPM) for direct learning signal distributions defined on the manifold mesh. We term our framework DoubleDiffusion, as it leverages heat dissipation diffusion to capture surface-specific features and employs the diffusion probabilistic framework for generative learning directly on the manifold. In summary, our contributions are:
  1. 1.
    We present DoubleDiffusion, the first framework for denoising diffusion probabilistic model that directly learns and generates signals on the 3D mesh surface, and achieves smoothness and view consistency along the mesh geometry.
  2. 2.
    Our framework achieves a significant 312.82% improvement in coverage compared with the previous SoTA MDF [12]. Meanwhile, it can process large meshes of size e.g., $\sim$100k vertices in a single pass,$\sim$20x more than MDF. It is also 8.1x faster to generate a sample on the manifold mesh.
  3. 3.
    Experiments on RGB distribution over a single manifold and on per-category shape-conditioned texture generation across multiple shapes demonstrate DoubleDiffusion’s versatility. These results highlight the framework’s capability for geometry-adaptive signal generation on complex 3D surfaces.
2 Related Work
2.1 Differential Geometry and Mesh Processing
In differential geometry, the surface of a 3D mesh is treated as a manifold, making learning over manifolds a longstanding topic of research in mesh processing. Taubinet al. [37] proposes using the Laplacian-Beltrami (LB) operator to define differential coordinates on meshes, noting that the integration of these coordinates over a small region approximates the mean curvature at a central point on the manifold. Later, Sörkineet al. [34] demonstrates that a mesh’s geometry can be effectively captured using the discrete Laplacian, enabling detailed geometric analysis on non-Euclidean surfaces. [43] indicates that the LB operator offers a robust approach for learning signals over mesh geometry. Inspired from this, our method follows the geometry processing theory and make use of the mesh Laplacian operators to process the signals over the manifold.
2.2 Diffusion on Manifold
Denoising Diffusion Probabilistic Models (DDPMs) are well-suited for learning on grid-like structured data, where neighborhood relationships are straightforwardly formulated with Euclidean distances. However, limited work has explored DDPMs within the context of manifold geometry. Manifold Diffusion Field (MDF) [12] and Diffusion Probabilistic Field (DPF) [46] address signal generation over the manifold through a field representation with the coordinate-signal pairs. In particular, MDF replaces Cartesian coordinates adopted in DPF with Laplacian eigenvectors to represent manifold geometry, and uses Transformer as the denoising network. Unlike these field-based approaches, our method directly processes signal distributions on manifold, integrating heat dissipation diffusion to capture local geometric dependencies and streamline the generative process.
2.3 Heat Diffusion
The heat equation has been widely applied in message propagation networks, particularly within graph-based deep learning. A closely related work by Xu et al. [41] leverages the heat kernel for semi-supervised learning on graphs, demonstrating that the Laplacian approximation over a heat kernel effectively filters out high-frequency components in node features, thereby enhancing signal smoothness. Similarly, Sharpet al. [31] introduces a heat dissipation diffusion network specifically for mesh structures, providing an ideal tool for our approach. In this work, we integrate the heat diffusion network [31] with a probabilistic diffusion generative model. Here, heat diffusion serves to suppress highly varying components in the noisy feature space, acting as an effective denoiser for the denoising diffusion probabilistic model on meshes.
2.4 Direct 3D Texture Generation
In 3D texture generation, challenges include the limited availability of high-quality 3D datasets and the high computational cost of training general diffusion models on mesh geometries. Some methods attempt to address these issues by projecting multiple views of the 3D object [45, 3] and applying 2D image diffusion models [44, 29] to synthesize signals on the mesh. However, these approaches often result in inconsistency in overlapping regions. Another set of approaches generates content on alternative 3D representations, such as Gaussian Splatting [36], Triplane [39], and Hash-grid [10], with subsequent mesh extraction, which frequently introduces artifacts and reduces quality [36, 10]. Differently, our work addresses these limitations by providing a framework for generating signals directly on 3D meshes in a way that respects the underlying geometry. This approach enables consistent, geometry-aware generation without the need for view-based projection or post-extraction process, thus preserving mesh integrity and fidelity.
3 Preliminary
3.1 Heat Diffusion
Heat diffusion has long been used in describing how an inital scalar heat distribution $\bm{h}$ diffuses or “spread out” over time ‘$s$’. In the differential geometry, it can describe how the signal diffuse over the continuous surfaces. Mathematically, this process is defined as a partial differential equation (PDE):
$\frac{\partial\bm{h}(s)}{\partial s}=\Delta\bm{h}(s),$ (1)
where $\bm{h}(s)$ represents the feature energy (or heat) over the manifold, $s$ represent the distance of how far the energy is transmitted, and $\Delta$ is the Laplacian Beltrami operator characterizes how the tempreature at a point dicerges from its surrounding region.
(a) Cotangent weights calculated on vertex $v_{i}$ w.r.t to edge (i,j)
(b) Cotangent of $\angle\alpha_{ij}$ is proportional to the area of triangle $F_{ijk}$, larger area in the left triangle will also result in a larger $\cot\alpha_{ij}$ comparing to the triangle in the right.
Figure 2: Illustration of the cotangent weighted differential coordinates on the manifold with respect to manifold mesh.
3.2 Mesh Laplacian
In the context of 3D mesh, that is, a discrete representation of surface manifolds composed of a finite number of vertices $\mathbf{V}$ and triangular faces $\mathbf{F}\subset\mathbf{V}^{3}$, we dicretize the continuous heat equation to operate on the mesh structure. And we use the cotangent-weighted Mesh Laplacian $L^{MC}$ proposed in [24] as a common practice.
$\mathbf{L}=M^{-1}\mathbf{C}.$ (2)
The cotangent weight matrix $C$ is calculated with respect to the angles adjacent to the shared edges, and the mass matrix is related to the area of the adjacent faces. These reflect the differential relationships and local geometry among vertices, acting similarly to an integration operation over the geodesic neighborhood around each vertex on manifold surfaces [4].
The cotangent weights $C$ are calculated as follows:
$C_{ij}=\sum_{j\in N(i)}\frac{\cot\alpha_{ij}+\cot\beta_{ji}}{2},$ (3)
where $C_{ij}\in\mathbb{R}$ represents the cotangent weight along the edge $(i,j)$, and $\alpha_{ij}$ and $\beta_{ji}$ are the angles opposite to edge $\bm{e}_{ij}$. This formulation considers both edge lengths and adjacent triangle areas (Fig.2), which are loosely related to the Gaussian curvature at the central vertex $v_{i}$, as stated in [24].
The cotangent is calculated by:
$\cot\alpha_{ij}$ $=\frac{\bm{e}_{ki}\cdot\bm{e}_{jk}}{|\bm{n}_{k}|},$ (4)
$\bm{e}_{ki}=\bm{v}_{i}-\bm{v}_{k},\bm{e}_{jk}$ $=\bm{v}_{j}-\bm{v}_{k},n_{k}=\bm{e}_{i}\times\bm{e}_{j},$
where $\bm{e}_{i}$ and $\bm{e}_{j}$ are the edge vectors.
When angles exceed $\frac{\pi}{2}$ , cotangent values become negative. To ensure a positive semi-definite Laplacian matrix, we follow common practice by applying a mass matrix $M\in\mathbb{R}^{n\times n}$, $M$ is a mass matrix given by: $M_{ii}=\frac{1}{3}\sum_{i\in t}\text{Area}(t),$ which assigns one-third of the area of each incident triangle to vertex $\bm{v}_{i}$, thus accounting for the vertex’s share of the surrounding area. Having the mass matrix into the cotan-weight Laplacian matrix makes it an FEM Laplacian operator and ensures that its eigenvalue decomposition is robust to mesh parameterization, as discussed in [43].
4 Method
4.1 Overview
                 (a)           (b)                      (c)                            (d)
Figure 3: Framework Overview. Our (a) DoubleDiffusion Network contains consecutive (b) DoubleDiffusion Blocks (DDB). Each DDB consists of the (c) Heat Diffusion module aggregating vertice features over the spatial domain on mesh surfaces and (d) the Timestep-Aware MLP that injects timestep embedding in a residual manner.
Our method begins by computing the cotangent-weighted Laplacian $\mathbf{L}$ from the manifold mesh structure (Sec.3.2), which serves as the foundation for diffusion on the manifold. Using this Laplacian, we perform eigen decomposition on the Laplacian-Beltrami operator, obtaining the smallest $k$ eigenvalues and eigenvectors.These are then used to approximate a heat diffusion kernel $\mathbf{H}(s)$across the manifold (Sec.4.2).
This heat dissipation diffusion acts as a geometric feature extractor on the manifold, where $s$ in $\mathbf{H}(s)$ represents the distance that a feature diffuses from vertex $\bm{v}_{i}$ across the manifold mesh $\mathcal{M}$. Analogous to a convolutional kernel on an image grid, this heat dissipation diffusion aggregates features on the scalar function $\bm{f}$ defined on the vertices of the manifold mesh as $\mathbf{L}\bm{f}$. Here, $\bm{f}=(\bm{f}^{v_{0}},\bm{f}^{v_{1}},\cdots,\bm{f}^{v_{n}})$, where $\bm{f}^{v_{1}}$ refers to the scalar vector on vertex $\bm{v}_{i}$, and $\{\bm{v}_{0},\bm{v_{1}},\cdots,\bm{v_{n}}\}\in\mathcal{M}$.
Our diffusion probabilistic model generates samples by starting with a noise distribution $\bm{f}_{T}^{v_{i}}\sim\mathcal{N}(0,1)$, the initial noise is independently sampled at each vertex. The DPM is trained to iteratively predict the reverse transitions $p(\bm{f}_{T}^{v_{i}}|\bm{f}_{T-1}^{v_{i}})$ progressively denoising through each step.
It’s worth noting that, we define two independent parameters: the heat dissipation diffusion distance $s$ and the denoising diffusion time step $t$ in the DoubleDiffusion framework. The diffusion distance $s$ is a learnable parameter within the heat dissipation network, allowing the model to adaptively capture geometric features on the manifold. In contrast, the denoising diffusion time step $t$ follows a predefined schedule, iteratively progressing from $T$ to 00 as part of a recursive Markov chain learning process.
4.2 Laplace Approximation of Heat Diffusion
Solving the heat equation presents an eigenvalue problem, where we can use the mesh Laplacian in Eq. 2 to approximate the heat diffusion. This approach also brings effect to fast compute to the heat diffusion Eq. 1 on large meshes. Meanwhile, Laplacian operators has been recognized as a spectral filer [37, 20]. Choosing the eigenvector of the mesh Laplacian $\mathbf{L}$ with $k$ smallest eigenvalues, Laplacian operator selectively attenuates high-frequency components (sharp changes or noise) while preserving low-frequency components (smooth, global variations). Thus, the Laplacian-approximated heat diffusion functions by dissipating high-frequency variations more quickly than low-frequency components. This behavior effectively smooths the signal, making the heat diffusion process an ideal tool for filtering.
We use the eigen-decomposition of the Laplacian operator in Eq. 2, yielding:
$L=\Phi\Lambda\Phi^{T},$ (5)
where $\Phi$ contains the eigenvectors of $L$, and $\Lambda$ is a diagonal matrix with eigenvalues $\lambda_{i}$. The heat diffusion is then approximated as:
$\mathbf{H}(s)\bm{f}=\Phi e^{-s\Lambda}\Phi^{T}\bm{f}=\sum_{i=1}^{k}e^{-s\lambda_{i}}\langle\bm{f},\phi_{i}\rangle\phi_{i}.$ (6)
In this spectral representation, the heat diffusion kernel $e^{-s\Lambda}$ has diagonal elements $e^{-s\lambda_{i}}$ that control the rate of decay for each frequency component.
Furthermore, we followed the method introduce in [31], to apply the heat kernal with respect to the feature channel:
$\mathbf{H}(s)\bm{f}=\Phi\begin{bmatrix}e^{-\lambda_{0}s}\\ e^{-\lambda_{1}s}\\ \vdots\end{bmatrix}\odot(\Phi^{T}\bm{f}),$ (7)
where the $\Phi$ only contains the $k$ smallest eigenvector of the $L^{MC}$ and the eigenvalues are in small to large order $\lambda_{0}<\lambda_{1}<\cdots<\lambda_{k}$. The symbol $\odot$ represents the element-wise multiplication. In this process, the $k$-eigenvectors first project the feature signals into the spectral domain, where the high-frequency components are suppressed by the frequency-modulated heat kernel, while low-frequency components are enhanced. This approach achieves channel-wise feature aggregation across vertices in the spectral domain, effectively gathering features within the heat diffusion range defined by the learnable $s$.
This connection between the heat diffusion equation and low-pass filtering allows us to effectively aggregate features across the manifold while preserving geometric consistency.
Denoising Process with Heat Diffusion. The reverse process $p_{\theta}(\bm{f}_{t-1}|\bm{f}_{t})$ removes noise iteratively. We parameterize this by predicting the noise $\bm{\epsilon}_{\theta}(\bm{f}_{t},t)$ using the heat diffusion kernel $\mathbf{H}(s)$, where $s$ is the heat diffusion distance:
$p_{\theta}(\bm{f}_{t-1}|\bm{f}_{t})=\mathcal{N}(\bm{f}_{t-1};\mu_{\theta}(\bm{f}_{t},t),\sigma_{t}^{2}\,\mathbf{I}),\$ (8)
where $\mu_{\theta}(\bm{f}_{t},t)=\frac{1}{\sqrt{1-\beta_{t}}}\left(\bm{f}_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\bm{\epsilon}_{\theta}(\bm{f}_{t},t)\right)$.
Following previous work [30], we incorporate the time step $t$ as a time embedding within each consecutive block using a per-vertex MLP. Given that heat diffusion is conducted on a per-channel basis, employing group normalization [40] directly disrupts the preservation of spatial information, resulting in non-convergence of the model. Consequently, we empirically adopt layer normalization [2] to enhance the training process.
Loss Function. We train by minimizing the MSE loss between the true noise $\bm{\epsilon}$ and predicted noise $\bm{\epsilon}_{\theta}(\bm{f}_{t},t)$:
$\mathcal{L}_{\text{DDPM}}=\mathbb{E}_{\bm{f}_{0},\bm{\epsilon},t}\left[\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\sqrt{\alpha_{t}}\,\bm{f}_{0}+\sqrt{1-\alpha_{t}}\,\bm{\epsilon},t)\right\|^{2}\right],$ (9)
where $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ and $t$ is sampled from $\{1,\dots,T\}$.
5 Experiments
In this section, we evaluate our proposed framework, DoubleDiffusion, for learning signal distributions on 3D mesh manifolds. We conduct three main experiments: (1) learning RGB color distributions on one prescribed manifold Stanford bunny (Sec. 5.1), (2) learning the texture distribution over a set of prescribe manifold and generates diverse textures conditioned on each shape’s geometry (Sec. 5.2). Both tasks are designed to test out the direct generative ability on the manifold as well as the potentials of the proposed method.
Figure 4: Qualitative comparison between the (a) MDF (baseline) and (b) DoubleDiffusion (ours) on the Stanford Bunny with $\sim$5k number of vertices.
Method MMD $\downarrow$ COV $\uparrow$ #param.$\downarrow$ $t$ (s) $\downarrow$
MDF 0.259 7.75 11.4M 1.14
Ours 0.284 32.01 2.4M 0.14
Table 1: Quantity comparison between MDF (baseline) and DoubleDiffusion (ours) with CelebA-HQ on Stanford Bunny (number of vertices $\sim$5k). With comparable MMD, our method surpasses MDF in terms of COV with a large margin. Meanwhile, our model is 4.7x smaller and 8.1x faster than MDF.
Figure 5: Our method can efficiently handle the mesh with large number of vertices manifold mesh. Here we show our result with (b) $\sim$25k vertices, and (c) $\sim$52k vertices (Stanford Bunny), with a qualitive comparison between (a) bunny with $\sim$5k vertices. The meshes with higher vertices are watertight remeshing with the off-the-shelf watertight manifold tool ManifoldPlus [16].
                   (a)                                                                                     (b)
Figure 6: Texture learning across multiple chair shapes in the ShapeNetCore v2  [7] dataset, with (a) the uncolored chair meshes and, (b) the textured chair meshes sampled from DoubleDiffusion.
5.1 Distribution Generation on Manifold
In this experiment, we evaluate the ability of our framework to learn complex RGB distributions on a 3D curved surfaces. The learning objective is to model the probability distribution $p_{\theta}(f^{v_{i}}_{t-1}|f^{v_{i}}_{t})$, where each $v_{i}\in\mathcal{M}$, with $\mathcal{M}$ representing the Stanford bunny manifold [38]. Here, $f^{v_{i}}_{t-1}$ is the vertex feature at denoising step $t-1$, with the initial feature $f_{t=0}^{v_{i}}$ being a 3-dimensional RGB value at each vertex acquired from Celeba-HQ human face dataset [18]. This setup allows us to assess the framework’s effectiveness in capturing a meaningful distribution over the manifold.
5.1.1 Implementation
Dataset Preparation. As there’s not an available datasets with a large number of texture for one manifold, we follow the baseline method MDF to prepare the manifold value data source from the CelebaA-HQ image, as the CelebA-HQ [18] dataset has complex and diverse RGB distributions. Initial RGB values $f_{t=0}^{v_{i}}$ are mapped onto the Stanford bunny manifold using an off-the-shelf texture mapping tool, PyVista [35], which provides a texture coordinate to the stanford bunny, and we then assigned the per-vertex color values from the image data in Celeba-HQ. Along with the manifold-color pair, we precompute the cotangent-weighted Laplacian matrix for the mesh, along with its eigenvalues and eigenvectors, to facilitate heat diffusion. The CelebA-HQ dataset consists of 30,000 images. We use the official split file with 24,183 images as the training set and 2,824 images as the test set.
Training Details. During training, the network predicts the denoising of RGB values across the manifold. The prediction network utilizes heat diffusion adapted from DiffusionNet [31] as a feature processing mechanism. We employ the EDM [19] framework as the denoising probabilistic model for faster inference with 18 timesteps. All experiments are conducted on 4 NVIDIA A100s, with 96 training epochs, learning rates at 3e-2, batch size of 8, 128 eigenvector to represent the approximation of the heat dissipation diffusion and 8 DoubleDiffusion blocks. The inference time is evaluated on the NVIDIA A100 GPU.
5.1.2 Evaluation
Baseline and Metrics. Due to the absence of official source code for MDF [12], we meticulously reproduce the baseline MDF method based on the detailed descriptions provided in their paper and its appendix. Specifically, we employed the Transf. Enc-Dec architecture as outlined in the appendix of MDF. We evaluated both MDF and our approach using a manifold-processed and simplified Stanford Bunny model, consisting of approximately 5k vertices [16]. This vertex count is significantly lower than the 50k+ vertices our model is capable of handling. The limitation is due to the computational constraints imposed by MDF’s global attention mechanism, which aggregates features across vertices and escalates computational complexity. For our evaluations, we adopted the metrics proposed by Achlioptas et al. [1]: Coverage (COV) and Minimum Matching Distance (MMD). These metrics assess the alignment between the test set and the generated samples, and indicate the accuracy with which the test set is represented within the generated samples, respectively. We followed the implementation from the official code111 https://github.com/optas/latent_3d_points/blob/master/notebooks/compute_evaluation_metrics.ipynb [1] to use the ground-truth and generated samples as reference and sample points, respectively, a practical contrary to that used by MDF. In particular, at equivalent levels of MMD a higher COV is desired [1], and vice-versa. To have a fair comparison, we generate equal number of data samples as the test set to compare the generated distribution with the test distribution.
Comparison. We first present a quantitative comparison in Tab. 1. Although our method posts comparable Minimum Matching Distance (MMD) scores, it achieves a significant 312.82% improvement in Coverage (COV), demonstrating superior alignment between the test-set and generated fields. Additionally, our model is 4.7 times smaller and 8.1 times faster than MDF, highlighting our efficient feature learning that leverages both heat diffusion and denoising diffusion processes. Fig. 4 showcases the qualitative comparison results. Our approach produces textures with enhanced details and more distinct facial features. In contrast, the textures generated by MDF appear excessively smoothed, resulting in a loss of fine details and a uniformly bland, less expressive surface appearance, which we refer to as ‘average faces’. This phenomenon aligns with the quantitative findings, where MDF tends to score better on MMD due to its tendency to predict average-like features thus fails to cover the test-set distribution. This discrepancy can be attributed to MDF’s reliance on Laplacian eigenmaps for spectral positional encodings and its formulation of texture generation as a field. During training, MDF must sample a limited number of points to stay within memory constraints, which may prevent it from fully capturing the underlying geometry. In contrast, our method facilitates local feature aggregation more attuned to the mesh’s shape, resulting in textures that more accurately reflect the surface’s curved structure.
5.1.3 Apply to Larger Meshes
We scaled our experiments to handle larger mesh with $24,992$ ($\sim$25k) and $52,299$ ($\sim$52k) vertices. Fig. 5 demonstrates our method can successfully generate comparatively high-quality textures on these higher-resolution meshes. With the support of more vertices, the generated samples capture better details with less noise. This scalability sheds lights on generative learning over large mesh with complex manifold surfaces, making DoubleDiffusion suitable for applications requiring detailed texture generation over the large number of vertices’ meshes, with a fast speed.
Note that it is infeasible for MDF to handle such high-resolution meshes. Facing the limitation of scalability, MDF can be only trained with a maximum of approximately 5,000 vertices on our default GPU. It is due to the adoption of the global attention mechanism to aggregate features at the vertices [25], which has quadratic time and memory complexity. While adopting the advanced framework like PerceiverIO [17] can handle this issue, the performance decreases with the reduce of the number of content points [12, 46]. Moreover, during training, MDF randomly samples points on the mesh surface to facilitate training, which can also hinder the modal to learn the underlying complete mesh geometry. In contrast, our method is specifically designed to learn signal distributions directly on the 3D manifolds by integrating heat dissipation diffusion with a probabilistic model. This combination allows to make a full use of the advantage of heat diffusion to effectively learn on complete mesh shape, and the ability of the denoising diffusion to generate high-quality signals.
5.2 Per-category Texture on ShapeNetCore
Figure 7: Generative result of the textured chairs from five distinct view points.
In this experiment, we extend DoubleDiffusion to learn over multiple manifold shapes. This setup models the distribution $p_{\theta}(f^{\mathcal{M}_{k,v_{i}}}_{t-1}|f^{\mathcal{M}_{k,v_{i}}}_{t})$, where $\mathcal{M}_{k,v_{i}}$ denotes vertex $v_{i}$ on manifold $\mathcal{M}_{k}$, with $\mathcal{M}_{k}\in\{\mathcal{M}_{m}\}$ representing a set of manifold shapes in the training data. This setup enables DoubleDiffusion to generate textures that are adapted to the structure of each individual mesh in the category, showcasing its capability to handle per-category texture generation across different manifolds.
5.2.1 Dataset and Implementation Details
We use the ShapeNetCore dataset [7] to perform a per-category texture generation. ShapeNetCore is a 3D objects dataset with 55 different categories, such as chairs, caps, cars, etc. Each shape object in the dataset have one corresponding texture. We use the area interpolation to acquire the texture color on the vertices of the shapes, as the texture atlas are defined in per-face manner. As a demonstration, we trained on the “chair” category, which includes 2,412 training samples and 311 test samples. We use the same network setting with the single mesh experiment (Sec. 5.1), only adapt the dataset to handle the batch shape with different number of vertices. To be able to handle shape with large number of vertices, we pad all the shape and related Laplacian matrix to 60,000. During training, we set the batchsize to 4, and train on 4 GPUs for 96 epochs, which is approximately 48 hours.
5.2.2 Qualitative results
Generative qualitive along the geometry. Figure 6 showcases qualitative results from our per-category texture generation experiment on the ShapeNetCore “chair” category. Each column represents a different chair model, where DoubleDiffusion generates textures that adapt to the unique geometric structure of each mesh. The leftmost images in each row show the untextured chair modelsFig. 6 demonstrate the qualitative result of texture generation on the different input chairs. The generated color on the mesh are smoothly varying following the chairs geometry, and the generated texture demonstrated a certain level of the understanding to the geometry.
In each example, the generated color on the mesh smoothly varies in alignment with the geometry of the chair, demonstrating the model’s capability to respect and adapt to surface details. Zoomed-in regions highlight the continuity and consistency of the generated textures, especially in curved and intricate parts of the mesh, such as the chair backs, armrests, and legs.
View Consistency. We present the generated results on ShapeNetCore models from five different views in Fig. 7. This multi-view presentation demonstrates DoubleDiffusion’s ability to achieve view consistency in signal generation directly on the mesh, without the need for 2D unrolling of the mesh. This consistency across views highlights the model’s capacity to respect the underlying mesh structure.
6 Limitation and Future Work
While the multi-view results demonstrate DoubleDiffusion’s strong capability for view-consistent texture generation directly on the mesh, it is also worth noting that our vertex-based approach relies heavily on vertex connectivity and the mesh Laplacian. This reliance can present challenges in accurately capturing color details on large faces, highlighting an area for future refinement in handling textures over varying face sizes on the mesh. Meanwhile, our experiment can only demonstrate the per-category shape conditioned texture generation, lacking the ability to generalize to texture generation on multi-class shapes. The limitation comes from the lack of enough shape-texture data pair, which could potentially be solved by leveraging 2D methods for texture supervision.
7 Conclusion
In this work, we introduced DoubleDiffusion, a novel framework that integrates heat dissipation diffusion with denoising diffusion models to directly generate geometry-respecting signal distributions on 3D meshes. Key contributions include a novel direct learning framework for diffusion on 3D manifold surfaces with the integration of heat dissipation diffusion for effective denoising aligned with mesh structure. We demonstrated that the propose framework performs geometry-adaptive signal generation across multiple complex surfaces.
Supplementary Material
In this supplementary material, we provide additional implementation details in  Appendix A, theoretical analysis in  Appendix B and qualitative results in Appendix C.
Appendix A Implementation Details
To give more precise description to the framework in Fig. 3, we have include the diagram to the detail implementation for data preparation in Fig. S1 and heat diffusion block in Fig. S2.
A.1 Input data preparation
To process the mesh structure, we begin by computing the cotangent-weighted Laplacian matrix as discussed in Sec. 4.1. Using eigenvalue decomposition, we then derive its eigenvalues and eigenvectors. Additionally, we compute the mass matrix and the $X$- and $Y$-basis vectors for the tangent plane at each vertex. These tangent plane bases can be utilized to scale the heat-diffused features via tangent gradients, following the approach outlined in [31].
Figure S1: Detail implementation of the input data preparation, for both mesh structure data and the vertex color data.
The extraction of vertex colors is crucial to our experiments, as no existing 3D dataset provides hundreds or thousands of color data points defined on a single mesh structure. Therefore, we follow the baseline experiments [12] to generate the color data distribution using the Stanford Bunny mesh. This process involves employing the off-the-shelf 3D visualization tool PyVista [35] to obtain texture coordinates for the vertices of the Stanford Bunny mesh. With these texture coordinates, we perform UV mapping to transfer image data from the CelebA-HQ dataset, resulting in a mapped color vector assigned to the vertices of the mesh.
For the ShapeNetCore dataset, we utilize PyTorch3D [ravi2020pytorch3d] to load the texture atlases corresponding to shapes in the chair category. In PyTorch3D, textures are represented as atlases associated with the mesh faces. To obtain vertex colors from these face-based textures, we perform weighted interpolation. Specifically, for each vertex, we identify all the triangular faces that include it. The area of each triangle is computed, and the texture colors of these triangles are weighted proportionally to their respective areas relative to the total area of all triangles containing the vertex. This ensures accurate color interpolation at the vertex level.
A.2 Implementation of Heat Diffusion Block
The computation of heat-diffused features on the mesh is illustrated in Fig. S2. This process takes place within the heat diffusion block, corresponding to sub-block (c) in Fig. 3. Here, the noised vertex color features are first weighted by the mass matrix to account for the areas of the surrounding faces. The weighted features are then transformed into the spectral domain using the eigenvectors of the mesh structure. In this domain, the features are scaled by heat diffusion coefficients, enabling adaptive learning of geometric features on the mesh.
Figure S2: Implementation details of the heat diffusion block
Next, the heat-diffused features are projected back to the original feature domain and concatenated with the noised color features to produce an aggregated feature. This aggregated feature is then passed to the Timestep-Aware Per-Vertex MLP, as depicted in block (d) of Fig. 3, to perform probabilistic denoised diffusion.
Appendix B Theoretical Analysis
In this section, we would like to provide more detail to justify why and how Laplacian Beltrami operator can be used for learning on the manifold surface.
B.1 Integral on the manifold
The intension of using the Laplacian Beltrami operator on the manifold and the mesh is arise from the laplacian’s integration formulation on the manifold surface, as advised in the Mesh Laplacian Processing [34].
To analyze and process scalar functions defined on a manifold, the Laplace-Beltrami operator provides a framework for differential and integral operations, analogous to how convolutional kernels operate in the image domain. Since a manifold is locally Euclidean, Sörkine [34] demonstrated that points on the manifold can be represented by differential coordinates $\delta$, which are conceptually similar to $\Delta x$ in one-dimensional Euclidean space:
$\delta_{i}$ $=\bm{v}_{i}-\frac{1}{d_{i}}\sum_{j\in N(i)}\bm{v}_{j}$ (S1)
$=\frac{1}{d_{i}}\sum_{j\in N(i)}(\bm{v}_{i}-\bm{v}_{j}),$
where $d_{i}$ is a weighted factor representing the degree of the vertex i, which is the number of connection of vertex i. $N(i)$ represent the neighboring vertex $\bm{v}_{j}$ of the vertex i, it can also be a cotangent weights as a common practice in manifold mesh processing (Sec 3.2). It’s worth noting that, if we take differential coordinates to the continuous form,
[…]
Eq.(S1) becomes an integration of vertex $\bm{v}_{i}$ among the infinite small region within the surface curve $\gamma$ with perimeter $|\gamma|$, and $\bm{v}$ are the points that lies in this infinite small closed surface curve. And the $\textit{Curv}(\bm{v}_{i})$ in right hand side of Eq.(S2) represent the mean curvature [34] at vertex $\bm{v}_{i}$, and $\bm{n}_{i}$ is the surface normal.
These differential coordinates can be fastly computed through laplacian beltrami operator as it is a discretization version of the integration operator on the manifold, where the differential geometry on the mesh is given by:
$L=I-D^{-1}A,$ (S3)
where $I\in\mathcal{R}^{n\times n}$ is the identity matrix, $D\in\mathcal{R}^{n\times n}$ is the diagonal matrix with the entries $D_{ii}$ representing the number of the connections of the neighboring vertices (degree). $A\in\mathcal{R}^{n\times n}$ is the adjacency matrix, $A_{ij}=1$ when the edge exits between vertex i and j.
By incorporating geodesic distance and the area of adjacent triangles, the differential coordinates on the manifold (Eq. S1) with respect to a central vertex can be reformulated using cotangent weights as:
$\delta_{ij}=\sum_{j\in N(i)}\frac{\cot\alpha_{ij}+\cot\beta_{ji}}{2}(\bm{v}_{i}-\bm{v}_{j})$ (S4)
where $\alpha_{ij}$ and $\beta_{ij}$ are the angles opposite the edge connecting vertices $i$ and $j$ in the triangles adjacent to this edge, these are illustrated in Fig. 2(a). Thus, the use of the cotangent weighted Laplacian matrix in Eq. 2 and Eq. 3 upon the manifold surface gives as the geometric properties of the mesh structure.
Appendix C Experiment Results
In this section, we provides additional experiment results for per-category shape conditioned texture generation result on other categories, such as trains Fig. S3, pianos Fig. S4 and faucets Fig. S5.
Figure S3: Additional qualitative result on train meshes.
Figure S4: Additional qualitative result on piano meshes.
Figure S5: Additional qualitative result on faucet meshes.